# Conformal Lyapunov Optimization: Optimal Resource Allocation under Deterministic Reliability Constraints

### v1

@@ VI Conclusions
This paper introduces conformal Lyapunov optimization (CLO), a novel optimization framework that addresses network cost minimization problems under strict, worst case and deterministic, reliability performance constraints. CLO integrates the standard optimization framework of Lyapunov optimization (LO), with the novel reliability mechanism of online conformal risk control. Simulation results have validated the theoretical guarantees of CLO in terms of long-term reliability performance, highlighting its advantages when compared with resource allocation strategies based on LO.
Future research directions may include the exploration of distributed implementations of CLO, as well as applications for settings involving multi-carrier transmissions, interfering users, and transmission outages.
Appendix A Proof of Proposition 1
Proof.
We firstly note that the long term reliability constraint (II-Ha) can be equivalently re-written as
$\frac{1}{F}\sum_{f=0}^{F-1}(\overline{L}_{f}^{k}-r^{k})\leq o(1).$ (39)
From O-CRC [12], we know that under Assumptions 1 and 2, and assuming that there exist constants $m,M$ such that $\theta^{k}_{f}\geq m$ and $\theta^{k}_{f}\leq M$, the update rule (30) yields the following chain of inequalities:
$\frac{m-\gamma-\theta^{k}_{0}}{F\gamma^{k}}\leq\frac{1}{F}\sum_{f=0}^{F-1}(\overline{L}_{f}^{k}-r^{k})\leq\frac{M+\gamma-\theta^{k}_{0}}{F\gamma^{k}},$ (40)
Looking at (40) it is clear how the left and the right-hand side of the chain tend to zero as the number of frames $F\to\infty$. Thus, according to the Squeeze Theorem, the constraint (II-Ha) is met. Considering a finite time horizon $T$, and recalling that $F=\frac{T}{S}$, it clear that the convergence speed increases as $S\to 1$, i.e., reducing the number of slots $S$ in each frame. ∎
@@ Appendix B Proof of Proposition 2
$\begin{split}\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\{J(t)\}&=\frac{1}{F}\sum_{f=0}^{F-1}\frac{1}{S}\sum_{t=fS+1}^{f(S+1)}\mathbb{E}\{J(t)\}\\ &\leq\frac{1}{F}\sum_{f=0}^{F-1}\left[J_{f}^{*}+\frac{\mathbb{E}\{G(fS+1)\}}{vs.}\right]+\frac{\mu}{V},\end{split}$ (41)
@@
where $\mu$ is a constant term [8]. Since we are assuming that $\mathbbm{E}\{G(fS+1)\}\leq\infty$, for a fixed value of the penalty parameter $V$, as $S\to\infty$, we end up with an approximate solution whose value is closer to the optimal value of the per-frame resource allocation problem $J_{f}^{*}$. Furthermore, by employing the upper-bound presented in Section I of the supplemental materials, and applying theorem 4.8 in [8] we also ensure that the LDPP function (cf. Section I of the supplementary items) is bounded for each slot $t\in[fS+1,f(S+1)]$ as follows
$\Delta_{p}(t)\leq k+VJ_{f}^{*}.$ (42)
@@
According to Theorem 4.2 in [8], this condition ensures the mean-rate stability of all the queues, as requested by constraint (II-Hb). ∎

### v5

@@ VI Conclusions
This paper introduces conformal Lyapunov optimization (CLO), a novel optimization framework that addresses optimal resource managements for network-based learning, under strict and deterministic constraints on the learning reliability. CLO integrates the standard optimization framework of Lyapunov optimization (LO), with the novel reliability mechanism of online conformal risk control. Simulation results have validated the theoretical guarantees of CLO in terms of long-term reliability performance, highlighting its advantages when compared with resource allocation strategies based on LO.
Future research directions may include the exploration of distributed implementations of CLO, as well as applications for more complex scenarios involving multi-carrier transmissions, interfering users, latency, and transmission outages.
Appendix A Monotonicity Proofs for Reliability and Precision Losses
Miscoverage and Set-Size Losses.
Let $\theta_{1}<\theta_{2}$, and define the prediction set as $\mathcal{C}(x,\theta)=\{y\in\mathcal{Y}:p(y|x)\geq\theta\}$. Since $p(y|x)\geq\theta_{2}$ implies $p(y|x)\geq\theta_{1}$, it follows that $\mathcal{C}(x,\theta_{2})\subseteq\mathcal{C}(x,\theta_{1})$. Consequently, $\mathbbm{1}(y_{\mathrm{true}}\notin\mathcal{C}(x,\theta_{1}))\leq\mathbbm{1}(y_{\mathrm{true}}\notin\mathcal{C}(x,\theta_{2}))$, showing that the miscoverage loss is non-decreasing with respect to parameter $\theta$.
For the set-size precision loss, since $\mathcal{C}(x,\theta_{2})\subseteq\mathcal{C}(x,\theta_{1})$, we have $|\mathcal{C}(x,\theta_{2})|\leq|\mathcal{C}(x,\theta_{1})|$. Dividing both sides by $|\mathcal{Y}|$ yields $\frac{|\mathcal{C}(x,\theta_{2})|}{|\mathcal{Y}|}\leq\frac{|\mathcal{C}(x,\theta_{1})|}{|\mathcal{Y}|}$, proving that the set-size precision loss is non-increasing with respect to parameter $\theta$. ∎
FNR and FPR losses.
Let $\theta_{1}<\theta_{2}$, and define the prediction set as $\mathcal{C}(x,\theta)=\{(i,j):p(i,j|x)\geq\theta\}$. Since $p(i,j|x)\geq\theta_{2}$ implies $p(i,j|x)\geq\theta_{1}$, it follows that $\mathcal{C}(x,\theta_{2})\subseteq\mathcal{C}(x,\theta_{1})$. For a fixed $y_{\mathrm{true}}$, we have $(y_{\mathrm{true}}\cap\overline{\mathcal{C}(x,\theta_{1})})\subseteq(y_{\mathrm{true}}\cap\overline{\mathcal{C}(x,\theta_{2})})$, and consequently, $\frac{|y_{\mathrm{true}}\cap\overline{\mathcal{C}(x,\theta_{1})}|}{|y_{\mathrm{true}|}}\leq\frac{|y_{\mathrm{true}}\cap\overline{\mathcal{C}(x,\theta_{2})}|}{|y_{\mathrm{true}}|}$, showing that the FNR loss is non-decreasing with respect to the parameter $\theta$.
For the FPR loss, since $\mathcal{C}(x,\theta_{2})\subseteq\mathcal{C}(x,\theta_{1})$, we have $(\overline{y}_{\mathrm{true}}\cap\mathcal{C}(x,\theta_{2}))\subseteq(\overline{y}_{\mathrm{true}}\cap\mathcal{C}(x,\theta_{1}))$, and hence $\frac{|\overline{y}_{\mathrm{true}}\cap\mathcal{C}(x,\theta_{2})|}{|\overline{y}_{\mathrm{true}}|}\leq\frac{|\overline{y}_{\mathrm{true}}\cap\mathcal{C}(x,\theta_{1})|}{|\overline{y}_{\mathrm{true}}|}$, proving that the FPR loss is non-increasing with respect to the parameter $\theta$. ∎
@@ Appendix B Proof of Proposition 1
Assuming the presence of a finite estimation and dissemination delay of frame loss information $d^{k}$, the long-term reliability loss at the $F$-th frame can be written as
$\frac{1}{F}\sum_{f=0}^{F-1}\overline{L}_{f}^{k}=\frac{1}{F}\left[\sum_{f=0}^{F-d^{k}-1}\overline{L}_{f}^{k}+\sum_{f=F-d^{k}}^{F-1}\overline{L}_{f}^{k}\right].$ (41)
The first sum in the right hand side of (41) can be written as
$(F-d^{k})\left[\frac{1}{F-d^{k}}\sum_{f=0}^{F-d^{k}-1}\overline{L}_{f}^{k}\right].$ (42)
Furthermore, from [12] the dynamic update of the reliability hyperparameters (30), leads to the following chain of inequalities
$r^{k}+\frac{m-\gamma^{k}-\theta_{0}^{k}}{(F-d^{k})\gamma^{k}}\leq\frac{1}{F-d^{k}}\sum_{f=0}^{F-d^{k}-1}\overline{L}_{f}^{k}\leq r^{k}+\frac{M+\gamma^{k}-\theta_{0}^{k}}{(F-d^{k})\gamma^{k}}.$ (43)
Thus, multiplying by $(F-d^{k})$, and taking into account that the second term in the right hand side of (41) is always in $[0,d^{k}]$ thanks to the boundedness assumption on the reliability loss, we end up with the following bounds for the average reliability loss at the $F$-th frame
$l(m)\leq\frac{1}{F}\sum_{f=0}^{F-1}\overline{L^{k}_{f}}\leq U(m),$ (44)
where the bounds are defined as $l(m)=r^{k}-\frac{r^{k}d^{k}}{F}+\frac{m-\gamma^{k}-\theta_{0}^{k}}{\gamma^{k}F}$, and $U(M)=r^{k}+\frac{M+\gamma^{k}-\theta_{0}^{k}}{\gamma^{k}F}+\frac{d^{k}(1-r^{k})}{F}$.
Appendix C Proof of Proposition 2
Proof.
@@ Appendix C Proof of Proposition 2
$\begin{split}\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\{J(t)\}&=\frac{1}{F}\sum_{f=0}^{F-1}\frac{1}{S}\sum_{t=fS+1}^{(f+1)S}\mathbb{E}\{J(t)\}\\ &\leq\frac{1}{F}\sum_{f=0}^{F-1}\left[J_{f}^{*}+\frac{\mathbb{E}\{G(fS+1)\}}{VS}\right]+\frac{\mu}{V},\end{split}$ (45)
@@
where $\mu$ is a constant term [8]. Since we are assuming that $\mathbbm{E}\{G(fS+1)\}\leq\infty$, for a fixed value of the penalty parameter $V$, as $S\to\infty$, we end up with an approximate solution whose value is closer to the optimal value of the per-frame resource allocation problem $J_{f}^{*}$. Furthermore, by employing the upper-bound presented in Section I of the supplemental materials, and applying theorem 4.8 in [8] we also ensure that the LDPP function (cf. Section I of the supplementary items) is bounded for each slot $t\in[fS+1,(f+1)S]$ as follows
$\Delta_{p}(t)\leq B+VJ_{f}^{*},$ (46)
@@
where $B$ is a constant term. According to Theorem 4.2 in [8], this condition ensures the mean-rate stability of all the queues, as requested by constraint (II-Hb). ∎
(a) FNR (23) as a function of the time slot index for users operating under different reliability constraints $(\eta=0.5)$.
(b) Energy vs. precision trade-off for users operating under different reliability constraints.
Figure 11: FNR evolution and average energy vs. precision trade-off for users operating under different long-term reliability constraints.
Appendix D Additional Results
D-A Heterogeneous Edge-Inference Scenarios
To demonstrate the applicability of the proposed approach to general scenarios, we evaluate the performance of CLO in remote inference scenarios where users operate under diverse reliability constraints. To this end, we consider $K=3$ EDs connected through the multi-hop network architecture depicted in Figure 5(b). The simulation parameters are consistent with those described in Section V-C, and the frame size is set to $S=10$ slots. We focus on a binary image segmentation task in which users operate under distinct long-term FNR constraints, namely $r^{k}=\{0.14,0.15,0.20\}$ for ED 1, ED 2, and ED 3, respectively.
Figure 11(a) depicts the evolution of the FNR over time for each user connected to the network. It can be observed that the long-term reliability of each user converges to the prescribed target value, thereby demonstrating the capacity of CLO to accommodate users operating under heterogeneous reliability requirements. Conversely, Figure 11(b) illustrates the trade-off between the average network energy consumption and the average precision, evaluated as $1-F_{s}(x,\theta)$, experienced by each device. Results have been averaged after $5,{000}$ time slots, at the convergence of the long-term reliability constraint. Due to the intrinsic trade-off between precision and reliability, the average precision degrades as the stringency of the long-term reliability constraint increases.
D-B Long-term Reliability under Finite Propagation and Estimation Delay
We test CLO over the multi-hop network architecture illustrated in Figure 5(b), using the same simulation parameters specified in Section V-C. We consider a frame size size $S=10$ slots, and uniform delay values $d^{k}=\{0,5,10\}$ frames are adopted for all users. Figure 12 analyzes the long-term reliability under varying delay conditions. As the delay $d^{k}$ increases, the violation of the long-term reliability constraint also grows. However, it consistently remains within the theoretical upper bounds, indicated by the dashed lines, which—according to Proposition 1—are within an additive error $\mathcal{O}(1/d^{k})$ from the ideal case (i.e., when $d^{k}=0$). This demonstrates that the signaling overhead due to the computation of the frame loss and the propagation of the updated reliability hyperparameters, do not prevent the algorithm from achieving the target reliability, but only affect the speed of convergence.
Figure 12: Long-term reliability under different estimation and propagation delay of the CLO hyperparameters update.
D-C Impact of the Precision Loss Approximation
To evaluate the performance degradation caused by imprecise precision loss estimation using low-complexity neural networks, we compare the proposed CLO strategy with a genie-aided resource allocation approach. In this benchmark, rather than using precision loss approximators, we directly utilize the segmentation networks deployed in the system to guide the resource allocation. While this approach lacks practical applicability, it serves as a meaningful upper bound, allowing us to quantify the degradation introduced by the use of low-complexity approximators in the resource allocation process.
To this aim, we assess the genie-aided resource allocation policy in the multi-hop network architecture reported in Figure 5(b). Figure 13 shows the average energy/precision trade-off reached by the two optimization strategies. The trade-off curves have been obtained simulating CLO for increasing values of the trade-off parameter $\eta$ in (26) over 10 independent realizations of the task sequence, and averaging the results over the last 1,000 slots, at the convergence of the reliability constraints. Figure 13 testifies that, on average, guiding the resource allocation policy with low-complexity neural network approximators leads to an overall performance loss around the range $1\%-2\%$.
Figure 13: Energy vs. precision trade-off for genie-aided CLO (solid lines) and for CLO driven by precision-loss approximators (dashed lines).